# A Unified View of Optimal Kernel Hypothesis Testing

### v1

@@ A.4 Proof sketch of pooled kernel separation
$\bigl{|}\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\leavevmode\nobreak\ S_{k}(\mathbb{X}_{N}^{\pi})-\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\leavevmode\nobreak\ S_{k}(\widetilde{\mathbb{X}}_{N}^{\pi})\bigr{|}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \Delta_{\mathrm{pool}}$ (93)
for any $\mathbb{X}_{N}$, $\widetilde{\mathbb{X}}_{N}$, $\pi$, as above. Adapting the reasoning of Equations 81, 89 and 84, we conclude that $\Delta_{\mathrm{mean}}\leq\Delta$, $\Delta_{\mathrm{max}}\leq\Delta$, and $\Delta_{\mathrm{fuse}}\leq\Delta$. In the (non-pooled) robust kernel separation setting of Section A.3, we observe that the quantity $r\Delta$ leads to the term $r/N$ in the kernel separation. In the pooled setting the quantity $\Delta_{\mathrm{pool}}$ leads to the same term $r/N$ for all three pooling mechanisms. We deduce that kernel pooling for robust tests leads to the uniform separation rate
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\leavevmode\nobreak\ \mathrm{Kdisc}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\biggl{\{}\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta),\log(|{\mathcal{K}}|)\}}{N}},\ \frac{r}{N}\biggr{\}}$ (94)
@@
$\max_{k\in{\mathcal{K}}}\,\mathrm{Kdisc}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\biggl{\{}\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta),\log(|{\mathcal{K}}|)\}}{N}},\ \frac{r}{N}\biggr{\}}$ (95)
@@ A.5 Proof sketch of efficient kernel separation
We detail the proof structure of the efficient kernel separation results: Equations 16, 31 and 47.
@@
where $U^{(1)},\dots,U^{(B)}$ are U-statistics on $\lfloor N/B\rfloor$ samples. The proof strategy follows the one of Section A.1 for the complete U-statistic, with the aim to relate $U_{\textrm{block}}$ to a non-negative $V_{\textrm{block}}=\frac{1}{B}\sum_{b=1}^{B}V^{(b)}$ with $V^{(b)}\coloneqq V\!\left(X_{1+(b-1)\lfloor N/B\rfloor},\dots,X_{b\lfloor N/B\rfloor}\right)$ for $b=1,\dots,B$. As in Equation 61, by Kim and Schrab (2023, Appendix E.13 and Lemma 22) for MMD and HSIC (and by direct computation for KSD), with sample size $\lfloor N/B\rfloor$, we have
$U^{(b)}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ V^{(b)}-\frac{2K}{\lfloor N/B\rfloor}$ (97)
@@
$U_{\textrm{block}}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ V_{\textrm{block}}-C_{1}\frac{B}{N}.$ (98)
As mentioned in Section 1.2, efficient tests using incomplete U-statistics with design ${\mathcal{D}}$ rely on the wild bootstrap to avoid having to compute new entries of the kernel/core matrix. The quantile obtained by wild bootstrap can be bounded with probability at least $21-\beta/$ 2 as
$q_{1-\alpha}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ {\frac{N}{|{\mathcal{D}}|}\log\biggl{(}\frac{1}{\alpha}\biggr{)}}$ (99)
@@
$q_{1-\alpha}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ C_{2}{\frac{B}{N}\log\biggl{(}\frac{1}{\alpha}\biggr{)}}$ (100)
with probability at least $21-\beta/$ 2.
The concentration inequalities for $\sqrt{V^{(1)}},\dots,\sqrt{V^{(B)}}$ of Equation 54 give
${\mathrm{Kdisc}}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sqrt{V^{(b)}}+C_{3}\sqrt{\frac{1}{N}\log\biggl{(}\frac{B}{\beta}\biggr{)}},$ (101)
which hold simultaneously for $b=1,\dots,B$ with probability at least $21-\beta/$ 2. Using Jensen’s inequality (finite form), we deduce that, with probability at least $21-\beta/$ 2, it holds that
${\mathrm{Kdisc}}\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ \frac{1}{B}\sum_{b=1}^{B}\sqrt{V^{(b)}}+C_{3}\sqrt{\frac{1}{N}\log\biggl{(}\frac{B}{\beta}\biggr{)}}$ (102)
$\leq\leavevmode\nobreak\ \sqrt{\frac{1}{B}\sum_{b=1}^{B}V^{(b)}}+C_{3}\sqrt{\frac{1}{N}\log\biggl{(}\frac{B}{\beta}\biggr{)}}$
$=\leavevmode\nobreak\ \sqrt{V_{\textrm{block}}}+C_{3}\sqrt{\frac{1}{N}\log\biggl{(}\frac{B}{\beta}\biggr{)}}$
$\leq\leavevmode\nobreak\ \sqrt{V_{\textrm{block}}}+C_{3}\sqrt{\frac{B}{N}\log\biggl{(}\frac{1}{\beta}\biggr{)}}.$
Then, similarly to the complete U-statistic reasoning of Equation 64, by combining Equations 98, 100 and 102, we get that
@@
${\mathrm{Kdisc}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{B\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}},$ (104)
@@ A.6 Proof sketch of pooled efficient kernel separation
The pooled efficient kernel separation results of Equations 17, 32 and 48 can simply be obtained by combining the reasoning of the pooled and efficient kernel separation results in Sections A.5 and A.4, respectively.
@@ A.7 Proof sketch of L2 separation
We detail the proof structure of the $L^{2}$ separation results: Equation 21 proved in Schrab et al. (2023, Corollary 7), Equation 36 proved in Schrab et al. (2022b, Theorem 3) (extension of Albert et al., 2022, Corollary 2, with theoretical quantiles, and of Kim et al., 2022, Proposition 8.7, with permuted quantiles), and Equation 49 proved here. Here, we improve the dependence in $\beta$ to be logarithmic for all these results. For L2 separation, we focus on the U-statistic case with $U_{0}\coloneqq U_{k}$ for some kernel $k$, with bootstrapped statistics ${U_{1}},\dots,{U_{B}}$.
@@ A.7.1 MMD two-sample testing
$\big{|}U_{0}-{\mathrm{Kdisc}}^{2}\big{|}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \sqrt{\frac{\sigma_{1}^{2}}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$ (105)
with probability at least $21-\beta/$ 2, where
$\sigma_{1}^{2}\leavevmode\nobreak\ \coloneqq\leavevmode\nobreak\ {\mathrm{var}_{Z}\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})\right]}\right)}\\ \leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ \left\lVert S_{\lambda}\psi\right\rVert_{L^{2}}^{2}$ (106)
@@
${\mathbb{P}}\!\left(\big{|}U_{b}\big{|}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{\tilde{C}_{1}}{N}\sqrt{\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}}\log\!\left(\frac{1}{\alpha}\right)\,\bigg{|}\,\mathbb{X}_{n},\mathbb{Y}_{n}\right)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \alpha.$ (107)
@@
$\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ \mathbb{E}\left[\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}\right]+\sqrt{\frac{\widetilde{\sigma}_{1}^{2}}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\leavevmode\nobreak\ \mathbb{E}\left[k_{\lambda}(Z,Z^{\prime})^{2}\right]+\sqrt{\frac{1}{N(\lambda_{1}\cdots\lambda_{d})^{2}}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\leavevmode\nobreak\ \frac{1}{\lambda_{1}\cdots\lambda_{d}}+\frac{1}{\lambda_{1}\cdots\lambda_{d}}\sqrt{\frac{1}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\leavevmode\nobreak\ \frac{1}{\lambda_{1}\cdots\lambda_{d}}\log\left(\frac{1}{\beta}\right)$ (108)
with probability at least $21-\beta/$ 2, where
$\widetilde{\sigma}_{1}^{2}\leavevmode\nobreak$ $\coloneqq\leavevmode\nobreak\ {\mathrm{var}_{Z}\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})^{2}\right]}\right)}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ {{\mathbb{E}}_{Z}\!\left[\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})^{2}\right]}\right)^{2}\right]}\leavevmode\nobreak\ \lesssim\leavevmode\nobreak\ {{\mathbb{E}}_{Z}\!\left[\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[k_{\lambda}(Z,Z^{\prime})^{2}\right]}\right)^{2}\right]}$ (109)
$\lesssim\leavevmode\nobreak\ {{\mathbb{E}}_{Z}\!\left[\!\left(\frac{1}{\lambda_{1}\cdots\lambda_{d}}\right)^{2}\right]}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{(\lambda_{1}\cdots\lambda_{d})^{2}}.$
@@
${q}_{1-\alpha}\leavevmode\nobreak$ $\lesssim\leavevmode\nobreak\ \frac{1}{N\sqrt{\lambda_{1}\cdots\lambda_{d}}}\sqrt{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right),$ (111)
with probability at least $21-\beta/$ 2.
@@
${\mathrm{Kdisc}}^{2}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}\left\lVert S_{\lambda}\psi\right\rVert_{L^{2}}^{2}+\frac{C}{2N\sqrt{\lambda_{1}\cdots\lambda_{d}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right)$ (113)
@@
The uniform separation of Equation 113 can then be expressed as
$\big{\|}\psi\big{\|}_{L^{2}}^{2}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \big{\|}\psi-S_{\lambda}\psi\big{\|}_{L^{2}}^{2}+\frac{C}{N\sqrt{\lambda_{1}\cdots\lambda_{d}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (115)
@@
$\big{\|}\psi-S_{\lambda}\psi\big{\|}_{L^{2}}^{2}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ c\big{\|}\psi\big{\|}_{L^{2}}^{2}+\widetilde{C}\sum_{i=1}^{d}\lambda^{2s}_{i}$ (116)
for some constants $c\in(0,1)$ and $\widetilde{C}>1$. Substituting this bound in the uniform separation condition of Equation 115, the type II error is guaranteed to be controlled by $\beta$ under the stronger requirement
$\big{\|}\psi\big{\|}_{L^{2}}^{2}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sum_{i=1}^{d}\lambda^{2s}_{i}+\frac{1}{N\sqrt{\lambda_{1}\cdots\lambda_{d}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (117)
@@
$\big{\|}\psi\big{\|}_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2s/(4s+d)}$ (118)

### v2

@@ A.4 Proof sketch of pooled kernel separation
$\bigl|\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\penalty 10000\ S_{k}(\mathbb{X}_{N}^{\pi})-\underset{k\in{\mathcal{K}}}{\mathrm{pool}}\penalty 10000\ S_{k}(\widetilde{\mathbb{X}}_{N}^{\pi})\bigr|\penalty 10000\ \leq\penalty 10000\ \Delta_{\mathrm{pool}}$ (93)
for any $\mathbb{X}_{N}$, $\widetilde{\mathbb{X}}_{N}$, $\pi$, as above. Adapting the reasoning of Equations˜81, 89 and 84, we conclude that $\Delta_{\mathrm{mean}}\leq\Delta$, $\Delta_{\mathrm{max}}\leq\Delta$, and $\Delta_{\mathrm{fuse}}\leq\Delta$. In the (non-pooled) robust kernel separation setting of Section˜A.3, we observe that the quantity $r\Delta$ leads to the term $r/N$ in the kernel separation. In the pooled setting the quantity $\Delta_{\mathrm{pool}}$ leads to the same term $r/N$ for all three pooling mechanisms. We deduce that kernel pooling for robust tests leads to the uniform separation rate
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\penalty 10000\ \mathrm{Kdisc}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \max\biggl\{\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta),\log(|{\mathcal{K}}|)\}}{N}},\ \frac{r}{N}\biggr\}$ (94)
@@
$\max_{k\in{\mathcal{K}}}\,\mathrm{Kdisc}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \max\biggl\{\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta),\log(|{\mathcal{K}}|)\}}{N}},\ \frac{r}{N}\biggr\}$ (95)
@@ A.5 Proof sketch of efficient kernel separation
We detail the proof structure of the efficient kernel separation results: Equations˜16, 31 and 47.
@@
where $U^{(1)},\dots,U^{(B)}$ are U-statistics on $\lfloor N/B\rfloor$ samples. The proof strategy follows the one of Section˜A.1 for the complete U-statistic, with the aim to relate $U_{\textrm{block}}$ to a non-negative $V_{\textrm{block}}=\frac{1}{B}\sum_{b=1}^{B}V^{(b)}$ with $V^{(b)}\coloneqq V\!\left(X_{1+(b-1)\lfloor N/B\rfloor},\dots,X_{b\lfloor N/B\rfloor}\right)$ for $b=1,\dots,B$. As in Equation˜61, by Kim and Schrab (2023, Appendix E.13 and Lemma 22) for MMD and HSIC (and by direct computation for KSD), with sample size $\lfloor N/B\rfloor$, we have
$U^{(b)}\penalty 10000\ \geq\penalty 10000\ V^{(b)}-\frac{2K}{\lfloor N/B\rfloor}$ (97)
@@
$U_{\textrm{block}}\penalty 10000\ \geq\penalty 10000\ V_{\textrm{block}}-C_{1}\frac{B}{N}.$ (98)
As mentioned in Section˜1.2, efficient tests using incomplete U-statistics with design $\mathcal{D}$} rely on the wild bootstrap to avoid having to compute new entries of the kernel/core matrix. The quantile obtained by wild bootstrap can be bounded with probability at least $1-\beta/2$ as
$q_{1-\alpha}\penalty 10000\ \lesssim\penalty 10000\ {\frac{N}{|{\mathcal{D}}|}\log\biggl(\frac{1}{\alpha}\biggr)}$ (99)
@@
$q_{1-\alpha}\penalty 10000\ \leq\penalty 10000\ C_{2}{\frac{B}{N}\log\biggl(\frac{1}{\alpha}\biggr)}$ (100)
with probability at least $1-\beta/2$.
The concentration inequalities for $\sqrt{V^{(1)}},\dots,\sqrt{V^{(B)}}$ of Equation˜54 give
${\mathrm{Kdisc}}\penalty 10000\ \leq\penalty 10000\ \sqrt{V^{(b)}}+C_{3}\sqrt{\frac{1}{N}\log\biggl(\frac{B}{\beta}\biggr)},$ (101)
which hold simultaneously for $b=1,\dots,B$ with probability at least $1-\beta/2$. Using Jensen’s inequality (finite form), we deduce that, with probability at least $1-\beta/2$, it holds that
[…]
$\leq\penalty 10000\ \sqrt{\frac{1}{B}\sum_{b=1}^{B}V^{(b)}}+C_{3}\sqrt{\frac{1}{N}\log\biggl(\frac{B}{\beta}\biggr)}$
$=\penalty 10000\ \sqrt{V_{\textrm{block}}}+C_{3}\sqrt{\frac{1}{N}\log\biggl(\frac{B}{\beta}\biggr)}$
$\leq\penalty 10000\ \sqrt{V_{\textrm{block}}}+C_{3}\sqrt{\frac{B}{N}\log\biggl(\frac{1}{\beta}\biggr)}.$
Then, similarly to the complete U-statistic reasoning of Equation˜64, by combining Equations˜98, 100 and 102, we get that
@@
${\mathrm{Kdisc}}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{B\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}},$ (104)
@@ A.6 Proof sketch of pooled efficient kernel separation
The pooled efficient kernel separation results of Equations˜17, 32 and 48 can simply be obtained by combining the reasoning of the pooled and efficient kernel separation results in Sections˜A.5 and A.4, respectively.
@@ A.7 Proof sketch of L2 separation
We detail the proof structure of the $L^{2}$ separation results: Equation˜21 proved in Schrab et al. (2023, Corollary 7), Equation˜36 proved in Schrab et al. (2022b, Theorem 3) (extension of Albert et al., 2022, Corollary 2, with theoretical quantiles, and of Kim et al., 2022, Proposition 8.7, with permuted quantiles), and Equation˜49 proved here. Here, we improve the dependence in $\beta$ to be logarithmic for all these results. For L2 separation, we focus on the U-statistic case with $U_{0}\coloneqq U_{k}$ for some kernel $k$, with bootstrapped statistics ${U_{1}},\dots,{U_{B}}$.
@@ A.7.1 MMD two-sample testing
$\big|U_{0}-{\mathrm{Kdisc}}^{2}\big|\penalty 10000\ \lesssim\penalty 10000\ \sqrt{\frac{\sigma_{1}^{2}}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$ (105)
with probability at least $1-\beta/2$, where
$\sigma_{1}^{2}\penalty 10000\ \coloneqq\penalty 10000\ {\mathrm{var}_{Z}\!\left({{\mathbb{E}}_{Z^{\prime}}\!\left[h_{\lambda}(Z,Z^{\prime})\right]}\right)}\\ \penalty 10000\ \lesssim\penalty 10000\ \left\lVert S_{\lambda}\psi\right\rVert_{L^{2}}^{2}$ (106)
@@
${\mathbb{P}}\!\left(\big|U_{b}\big|\penalty 10000\ \geq\penalty 10000\ \frac{\tilde{C}_{1}}{N}\sqrt{\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}}\log\!\left(\frac{1}{\alpha}\right)\,\bigg|\,\mathbb{X}_{n},\mathbb{Y}_{n}\right)\penalty 10000\ \leq\penalty 10000\ \alpha.$ (107)
@@
$\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}\penalty 10000\$$\leq\penalty 10000\ \mathbb{E}\left[\frac{1}{N(N-1)}\sum_{1\leq i\neq j\leq N}h_{\lambda}(Z_{i},Z_{j})^{2}\right]+\sqrt{\frac{\widetilde{\sigma}_{1}^{2}}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\penalty 10000\ \mathbb{E}\left[k_{\lambda}(Z,Z^{\prime})^{2}\right]+\sqrt{\frac{1}{N(\lambda_{1}\cdots\lambda_{d})^{2}}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\penalty 10000\ \frac{1}{\lambda_{1}\cdots\lambda_{d}}+\frac{1}{\lambda_{1}\cdots\lambda_{d}}\sqrt{\frac{1}{N}\log\left(\frac{1}{\beta}\right)}+\frac{1}{N}\log\left(\frac{1}{\beta}\right)$
$\lesssim\penalty 10000\ \frac{1}{\lambda_{1}\cdots\lambda_{d}}\log\left(\frac{1}{\beta}\right)$ (108)
with probability at least $1-\beta/2$, where
[…]
$\lesssim\penalty 10000\ {{\mathbb{E}}_{Z}\!\left[\!\left(\frac{1}{\lambda_{1}\cdots\lambda_{d}}\right)^{2}\right]}\penalty 10000\ =\penalty 10000\ \frac{1}{(\lambda_{1}\cdots\lambda_{d})^{2}}.$
@@
${q}_{1-\alpha}\penalty 10000\$$\lesssim\penalty 10000\ \frac{1}{N\sqrt{\lambda_{1}\cdots\lambda_{d}}}\sqrt{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right),$ (111)
with probability at least $1-\beta/2$.
@@
${\mathrm{Kdisc}}^{2}\penalty 10000\ \geq\penalty 10000\ \frac{1}{2}\left\lVert S_{\lambda}\psi\right\rVert_{L^{2}}^{2}+\frac{C}{2N\sqrt{\lambda_{1}\cdots\lambda_{d}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right)$ (113)
@@
The uniform separation of Equation˜113 can then be expressed as
$\big\|\psi\big\|_{L^{2}}^{2}\penalty 10000\ \geq\penalty 10000\ \big\|\psi-S_{\lambda}\psi\big\|_{L^{2}}^{2}+\frac{C}{N\sqrt{\lambda_{1}\cdots\lambda_{d}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (115)
@@
$\big\|\psi-S_{\lambda}\psi\big\|_{L^{2}}^{2}\penalty 10000\ \leq\penalty 10000\ c\big\|\psi\big\|_{L^{2}}^{2}+\widetilde{C}\sum_{i=1}^{d}\lambda^{2s}_{i}$ (116)
for some constants $c\in(0,1)$ and $\widetilde{C}>1$. Substituting this bound in the uniform separation condition of Equation˜115, the type II error is guaranteed to be controlled by $\beta$ under the stronger requirement
$\big\|\psi\big\|_{L^{2}}^{2}\penalty 10000\ \gtrsim\penalty 10000\ \sum_{i=1}^{d}\lambda^{2s}_{i}+\frac{1}{N\sqrt{\lambda_{1}\cdots\lambda_{d}}}{\log\left(\frac{1}{\beta}\right)}\log\!\left(\frac{1}{\alpha}\right).$ (117)
@@
$\big\|\psi\big\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2s/(4s+d)}$ (118)